# Quantum Droplets of Light in Semiconductor Microcavities

### v1

@@ I.1 Interaction strength and polariton T-matrix
|$\ket{L_{\sigma},L_{\sigma^{\prime}},\mathbf{p}}=\hat{L}_{-\mathbf{p}\sigma}^{\dagger}\hat{L}_{\mathbf{p}\sigma^{\prime}}^{\dagger}\ket{0}.$ (S3)
@@
$\bra{L_{\sigma},L_{\sigma^{\prime}},\mathbf{p}^{\prime}}\hat{V}\ket{L_{\sigma},L_{\sigma^{\prime}},\mathbf{p}}$ $=(1+\delta_{\sigma\sigma^{\prime}})|X_{\mathbf{p}}|^{4}v_{\sigma\sigma^{\prime}}$ (S4a)
$\bra{L_{\sigma},L_{\sigma^{\prime}},\mathbf{p}^{\prime}}\hat{V}\frac{1}{E-\hat{H}_{0}}\hat{V}\ket{L_{\sigma},L_{\sigma^{\prime}},\mathbf{p}}$ $=(1+\delta_{\sigma\sigma^{\prime}})|X_{\mathbf{p}}|^{4}v_{\sigma\sigma^{\prime}}^{2}\Pi(E),$ (S4b)
@@
$T_{\sigma\sigma^{\prime}}(p)$ $\equiv\frac{\bra{L_{\sigma},L_{\sigma^{\prime}},\mathbf{p}^{\prime}}\hat{T}(2E_{\rm LP}(\mathbf{p}))\ket{L_{\sigma},L_{\sigma^{\prime}},\mathbf{p}}}{1+\delta_{\sigma\sigma^{\prime}}}$
@@ I.2 Bogoliubov approximation
$\Phi=-\frac{v_{\uparrow\downarrow}}{\mathcal{A}}\sum_{\textbf{p}}\langle\hat{x}_{\textbf{p}\uparrow}\hat{x}_{-\textbf{p}\downarrow}\rangle$ (S15)
@@ I.4 LHY quantum fluctuations
~$\tilde{\mathcal{H}}(\mathbf{k})=\begin{pmatrix}\epsilon_{\mathbf{k}}^{\rm C}-\mu_{\uparrow}&0&0&0&\Omega_{\rm R}/2&0&0&0\\ 0&\epsilon_{\mathbf{k}}^{\rm C}-\mu_{\downarrow}&0&0&0&\Omega_{\rm R}/2&0&0\\ 0&0&-\epsilon_{\mathbf{k}}^{\rm C}+\mu_{\uparrow}&0&0&0&\Omega_{\rm R}/2&0\\ 0&0&0&-\epsilon_{\mathbf{k}}^{\rm C}+\mu_{\downarrow}&0&0&0&\Omega_{\rm R}/2\\ \Omega_{\rm R}/2&0&0&0&\epsilon^{\rm X}_{\mathbf{k}}-\mu_{\uparrow}+2gn_{0\uparrow}&0&gn_{0\uparrow}&-\Phi\\ 0&\Omega_{\rm R}/2&0&0&0&\epsilon^{\rm X}_{\mathbf{k}}-\mu_{\downarrow}+2gn_{0\downarrow}&-\Phi&gn_{0\downarrow}\\ 0&0&\Omega_{\rm R}/2&0&-gn_{0\uparrow}&\Phi^{*}&-\epsilon^{\rm X}_{\mathbf{k}}+\mu_{\uparrow}-2gn_{0\uparrow}&0\\ 0&0&0&\Omega_{\rm R}/2&\Phi^{*}&-gn_{0\downarrow}&0&-\epsilon^{\rm X}_{\mathbf{k}}+\mu_{\downarrow}-2gn_{0\downarrow}\end{pmatrix},$ (S23)
which has the eight eigenvalues ±$\pm E^{a}_{\pm,\mathbf{k}}$, ±$\pm E^{b}_{\pm,\mathbf{k}}$. The Hamiltonian matrix $\mathcal{H}(\mathbf{k})$ is diagonalized by the Bogoliubov transformation $\hat{\Psi}_{\mathbf{k}}=\mathcal{R}\hat{\phi}_{\mathbf{k}}$, with $\hat{\phi}_{\mathbf{k}}=\begin{pmatrix}\hat{a}_{\mathbf{k},+}&\hat{a}_{\mathbf{k},-}&\hat{a}_{-\mathbf{k},+}^{\dagger}&\hat{a}_{-\mathbf{k},-}^{\dagger}&\hat{b}_{\mathbf{k},+}&\hat{b}_{\mathbf{k},-}&\hat{b}_{-\mathbf{k},+}^{\dagger}&\hat{b}_{-\mathbf{k},-}^{\dagger}\end{pmatrix}^{T}$. The columns of the matrix $\mathcal{R}$ are (properly normalized) eigenvectors of ~$\tilde{\mathcal{H}}(\mathbf{k})$ corresponding, respectively,to the eigenvalues $E^{a}_{+,\mathbf{k}}$, $E^{a}_{-,\mathbf{k}}$, $-E^{a}_{+,\mathbf{k}}$, $-E^{a}_{-,\mathbf{k}}$, $E^{b}_{+,\mathbf{k}}$, $E^{b}_{-,\mathbf{k}}$, $-E^{b}_{+,\mathbf{k}}$ and $-E^{b}_{-,\mathbf{k}}$. Applying this transformation, one obtains
$\hat{\Psi}_{\mathbf{k}}^{\dagger}\mathcal{R}\,\mathcal{R}^{\dagger}\mathcal{H}(\mathbf{k})\mathcal{R}\,\mathcal{R}^{\dagger}\hat{\Psi}_{\mathbf{k}}$
$=\begin{pmatrix}\hat{a}_{\mathbf{k},+}^{\dagger}&\hat{a}_{\mathbf{k},-}^{\dagger}&\hat{a}_{-\mathbf{k},+}&\hat{a}_{-\mathbf{k},-}&\hat{b}_{\mathbf{k},+}^{\dagger}&\hat{b}_{\mathbf{k},-}^{\dagger}&\hat{b}_{-\mathbf{k},+}&\hat{b}_{-\mathbf{k},-}\end{pmatrix}\begin{pmatrix}E_{+,\mathbf{k}}^{a}&&&&&&&\\ &E_{-,\mathbf{k}}^{a}&&&&&&\\ &&E_{+,\mathbf{k}}^{a}&&&&&\\ &&&E_{-,\mathbf{k}}^{a}&&&&\\ &&&&E_{+,\mathbf{k}}^{b}&&&\\ &&&&&E_{-,\mathbf{k}}^{b}&&\\ &&&&&&E_{+,\mathbf{k}}^{b}&\\ &&&&&&&E_{-,\mathbf{k}}^{b}\end{pmatrix}\begin{pmatrix}\hat{a}_{\mathbf{k},+}\\ \hat{a}_{\mathbf{k},-}\\ \hat{a}_{-\mathbf{k},+}^{\dagger}\\ \hat{a}_{-\mathbf{k},-}^{\dagger}\\ \hat{b}_{\mathbf{k},+}\\ \hat{b}_{\mathbf{k},-}\\ \hat{b}_{-\mathbf{k},+}^{\dagger}\\ \hat{b}_{-\mathbf{k},-}^{\dagger}\end{pmatrix}$
@@ II Symmetric spin mixture
$\Omega_{\rm MF}(\alpha,\Phi,\mu)=-\frac{\Phi^{2}}{g_{\uparrow\downarrow}}-\left(\Phi+\mu+\frac{\Omega_{\rm R}^{2}}{4\left(\delta_{0}-\mu\right)}\right)n_{0}+\frac{g}{4}n_{0}^{2},$ (S26)
@@
${E_{\pm,\mathbf{k}}^{a}}^{2}$ $=\frac{A_{\mathbf{k}}^{2}+B_{\mathbf{k}}^{2}-C_{+}^{2}}{2}+\left(\frac{\Omega_{\rm R}}{2}\right)^{2}\pm\sqrt{\left(\frac{A_{\mathbf{k}}^{2}-B_{\mathbf{k}}^{2}+C_{+}^{2}}{2}\right)^{2}+\left[\left(A_{\mathbf{k}}+B_{\mathbf{k}}\right)^{2}-C_{+}^{2}\right]\left(\frac{\Omega_{\rm R}}{2}\right)^{2}}$ (S30)
@@
0 0 $=\left[A_{0}(B_{0}-C_{+})-\frac{\Omega_{\rm R}^{2}}{4}\right]\left[A_{0}(B_{0}+C_{+})-\frac{\Omega_{\rm R}^{2}}{4}\right]$
@@
    $E_{\pm,\mathbf{k}}^{a}\,$ $\mapsto\,E_{-}(\mathbf{k})=\sqrt{\epsilon_{\mathbf{k}}^{\rm X}\left(\epsilon_{\mathbf{k}}^{\rm X}+gn_{0}+2\Phi\right)}$
@@ II.1 Limit of small photon mass
[…]
@@
[…]
@@ III Green’s function
where $C_{\sigma}(\mathbf{r})$/$X_{\sigma}(\mathbf{r})$ are the annihiliation field operators of a photon/exciton of species $\sigma=\uparrow,\downarrow$. Within a general formalism at finite temperature ($\beta=1/k_{\rm B}T$), the action of the system is given by:
[…]
@@
[…]
@@
[…]
In order to analyze the gaussian kernel, it is convenient to perform a Fourier transform and work in momentum space $(i\omega_{m},\mathbf{k})$, where $\omega_{m}=2\pi m/\beta$ ($m\in\mathbb{Z}$) is a bosonic Matsubara frequency. The results at zero temperature can then be obtained through the analytic continuation $-\hbar\partial_{\tau}\rightarrow i\omega_{m}\rightarrow\omega+i0^{+}$. The (inverse) Green’s function of the system is given by
@@
where we have restricted to the symmetric mixtures of lower polaritons that we analyze in the main text. Notice that the poles of the Green’s function, given by the condition $\text{det}\left[\mathcal{G}^{-1}(\mathbf{k},\omega\rightarrow E_{\mathbf{k}})\right]=0$, coincide with the eigenvalues of the Bogoliubov matrix ~$\tilde{\mathcal{H}}(\mathbf{k})$ defined in Eq. (S23).
@@ V.2 Spatial density profile in the slab-geometry configuration
[…]
@@
    numerically solve the second order ODE (S63) using the initial conditions from i) in the interval $[-x_{*},+x_{*}]$, where $x_{*}>0$ is such that $n_{0}(-x_{*})=n_{0\rm D}$;

### v3

@@ I Bogoliubov theory with bosonic pairing
Formally, our system is reminiscent of the scenario where two bosonic species of atoms are Rabi coupled, which has also been shown to feature droplet formation Cappellaro et al. (2017); Lavoine et al. (2021); Chiquillo (2025). However, we emphasize that there are several crucial differences. In our work, the two matter components are coupled separately to two photon components, rather than being coupled to each other. This leads to a system with four distinct components (two matter and two light, with two Rabi couplings), rather than two components with a single Rabi coupling. Furthermore, the light-matter coupling fundamentally alters the system as, in the absence of many-body effects, it leads to new part-light, part-matter exciton-polariton quasiparticles with a non-parabolic dispersion, something that cannot occur in ultracold atomic gases.
@@ I.1 Interaction strength and polariton T-matrix
[…]
@@
[…]
@@
[…]
@@ I.2 Bogoliubov approximation
$\Phi=-\frac{v_{\uparrow\downarrow}}{\mathcal{A}}\sum_{\textbf{p}}\langle\hat{x}_{\textbf{p}\uparrow}\hat{x}_{-\textbf{p}\downarrow}\rangle,$ (S15)
@@ I.4 LHY quantum fluctuations
$\tilde{\mathcal{H}}(\mathbf{k})=\begin{pmatrix}\epsilon_{\mathbf{k}}^{\rm C}-\mu_{\uparrow}&0&0&0&\Omega_{\rm R}/2&0&0&0\\ 0&\epsilon_{\mathbf{k}}^{\rm C}-\mu_{\downarrow}&0&0&0&\Omega_{\rm R}/2&0&0\\ 0&0&-\epsilon_{\mathbf{k}}^{\rm C}+\mu_{\uparrow}&0&0&0&-\Omega_{\rm R}/2&0\\ 0&0&0&-\epsilon_{\mathbf{k}}^{\rm C}+\mu_{\downarrow}&0&0&0&-\Omega_{\rm R}/2\\ \Omega_{\rm R}/2&0&0&0&\epsilon^{\rm X}_{\mathbf{k}}-\mu_{\uparrow}+2gn_{0\uparrow}&0&gn_{0\uparrow}&-\Phi\\ 0&\Omega_{\rm R}/2&0&0&0&\epsilon^{\rm X}_{\mathbf{k}}-\mu_{\downarrow}+2gn_{0\downarrow}&-\Phi&gn_{0\downarrow}\\ 0&0&-\Omega_{\rm R}/2&0&-gn_{0\uparrow}&\Phi^{*}&-\epsilon^{\rm X}_{\mathbf{k}}+\mu_{\uparrow}-2gn_{0\uparrow}&0\\ 0&0&0&-\Omega_{\rm R}/2&\Phi^{*}&-gn_{0\downarrow}&0&-\epsilon^{\rm X}_{\mathbf{k}}+\mu_{\downarrow}-2gn_{0\downarrow}\end{pmatrix}$, (S23)
which has the eight eigenvalues $\pm E^{a}_{\pm,\mathbf{k}}$, $\pm E^{b}_{\pm,\mathbf{k}}$. The Hamiltonian matrix $\mathcal{H}(\mathbf{k})$ is diagonalized by the Bogoliubov transformation $\hat{\Psi}_{\mathbf{k}}=\mathcal{R}\hat{\phi}_{\mathbf{k}}$, with $\hat{\phi}_{\mathbf{k}}=\begin{pmatrix}\hat{a}_{\mathbf{k},+}&\hat{a}_{\mathbf{k},-}&\hat{b}_{\mathbf{k},+}&\hat{b}_{\mathbf{k},-}&\hat{a}_{-\mathbf{k},+}^{\dagger}&\hat{a}_{-\mathbf{k},-}^{\dagger}&\hat{b}_{-\mathbf{k},+}^{\dagger}&\hat{b}_{-\mathbf{k},-}^{\dagger}\end{pmatrix}^{T}$. The columns of the matrix $\mathcal{R}$ are (properly normalized) eigenvectors of $\tilde{\mathcal{H}}(\mathbf{k})$ corresponding, respectively, to the eigenvalues $E^{a}_{+,\mathbf{k}}$, $E^{a}_{-,\mathbf{k}}$, $E^{b}_{+,\mathbf{k}}$, $E^{b}_{-,\mathbf{k}}$, $E^{a}_{+,\mathbf{k}}$, $E^{a}_{-,\mathbf{k}}$, $E^{b}_{+,\mathbf{k}}$ and $E^{b}_{-,\mathbf{k}}$. Applying this transformation, one obtains
$\hat{\Psi}_{\mathbf{k}}^{\dagger}\left(\mathcal{R}^{\dagger}\right)^{-1}\,\mathcal{R}^{\dagger}\mathcal{H}(\mathbf{k})\mathcal{R}\,\mathcal{R}^{-1}\hat{\Psi}_{\mathbf{k}}$
$=\begin{pmatrix}\hat{a}_{\mathbf{k},+}^{\dagger}&\hat{a}_{\mathbf{k},-}^{\dagger}&\hat{b}_{\mathbf{k},+}^{\dagger}&\hat{b}_{\mathbf{k},-}^{\dagger}&\hat{a}_{-\mathbf{k},+}&\hat{a}_{-\mathbf{k},-}&\hat{b}_{-\mathbf{k},+}&\hat{b}_{-\mathbf{k},-}\end{pmatrix}\begin{pmatrix}E_{+,\mathbf{k}}^{a}&&&&&&&\\ &E_{-,\mathbf{k}}^{a}&&&&&&\\ &&E_{+,\mathbf{k}}^{b}&&&&&\\ &&&E_{-,\mathbf{k}}^{b}&&&&\\ &&&&E_{+,\mathbf{k}}^{a}&&&\\ &&&&&E_{-,\mathbf{k}}^{a}&&\\ &&&&&&E_{+,\mathbf{k}}^{b}&\\ &&&&&&&E_{-,\mathbf{k}}^{b}\end{pmatrix}\begin{pmatrix}\hat{a}_{\mathbf{k},+}\\ \hat{a}_{\mathbf{k},-}\\ \hat{b}_{\mathbf{k},+}\\ \hat{b}_{\mathbf{k},-}\\ \hat{a}_{-\mathbf{k},+}^{\dagger}\\ \hat{a}_{-\mathbf{k},-}^{\dagger}\\ \hat{b}_{-\mathbf{k},+}^{\dagger}\\ \hat{b}_{-\mathbf{k},-}^{\dagger}\end{pmatrix}$
@@ II Symmetric spin mixture
$\Omega_{\rm MF}(\alpha,\Phi,\mu)=-\frac{\Phi^{2}}{g_{\uparrow\downarrow}}-\left(\Phi+\mu+\frac{\Omega_{\rm R}^{2}}{4\left(\delta_{0}-\mu\right)}\right)n_{0}+\frac{g}{4}n_{0}^{2}\,.$ (S26)
@@
${E_{\pm,\mathbf{k}}^{a}}^{2}$ $=\frac{A_{\mathbf{k}}^{2}+B_{\mathbf{k}}^{2}-C_{+}^{2}}{2}+\left(\frac{\Omega_{\rm R}}{2}\right)^{2}\pm\sqrt{\left(\frac{A_{\mathbf{k}}^{2}-B_{\mathbf{k}}^{2}+C_{+}^{2}}{2}\right)^{2}+\left[\left(A_{\mathbf{k}}+B_{\mathbf{k}}\right)^{2}-C_{+}^{2}\right]\left(\frac{\Omega_{\rm R}}{2}\right)^{2}}\,$, (S30)
@@
$0$ $=\left[A_{0}(B_{0}-C_{+})-\frac{\Omega_{\rm R}^{2}}{4}\right]\left[A_{0}(B_{0}+C_{+})-\frac{\Omega_{\rm R}^{2}}{4}\right]$
@@
    $E_{\pm,\mathbf{k}}^{a}\,$ $\mapsto\,E_{-}(\mathbf{k})=\sqrt{\epsilon_{\mathbf{k}}^{\rm X}\left(\epsilon_{\mathbf{k}}^{\rm X}+gn_{0}+2\Phi\right)}\,$,
@@ II.1 Limit of small photon mass
$\frac{1}{2\pi}\int_{0}^{\infty}dk\,k\,\left(\sqrt{B_{\mathbf{k}}^{2}-C_{s}^{2}}-B_{\mathbf{k}}+\frac{C_{s}^{2}}{2B_{\mathbf{k}}}\right)$
$=\frac{m_{\rm X}C_{s}^{2}}{2\pi\hbar^{2}}\,\int_{B_{0}/C_{s}-1}^{\infty}dy\left[\sqrt{y(y+2)}-(y+1)+\frac{1}{2y+2}\right]$
@@
$\frac{C_{s}^{2}}{2\pi}\int_{0}^{\infty}dk\left(\frac{k}{\hbar^{2}k^{2}/m_{\rm X}+2|\epsilon_{\rm LP}|}-\frac{k}{2B_{\mathbf{k}}}\right)=\frac{m_{\rm X}C_{s}^{2}}{4\pi\hbar^{2}}\ln\left(\frac{B_{0}}{|\epsilon_{\rm LP}|}\right).$ (S37)
@@ III Green’s function
where $C_{\sigma}(\mathbf{r})$$X_{\sigma}(\mathbf{r})$ are the annihiliation field operators of a photon/exciton of species $\sigma=\uparrow,\downarrow$. Within a general formalism at finite temperature ($\beta=1/k_{\rm B}T$), the action of the system is given by:
$\mathcal{S}=$ $\int d\mathbf{r}\int_{0}^{\beta\hbar}d\tau\,\left\{\sum_{\sigma}\left[C_{\sigma}^{\dagger}(\mathbf{r},\tau)\hbar\partial_{\tau}C_{\sigma}(\mathbf{r},\tau)+X_{\sigma}^{\dagger}(\mathbf{r},\tau)\hbar\partial_{\tau}X_{\sigma}(\mathbf{r},\tau)\right]+\mathscr{H}(\mathbf{r},\tau)\right\}.$ (S40)
@@
which allows to decompose the total action into $\mathcal{S}=\mathcal{S}_{\rm MF}+\mathcal{S}_{\rm B}.$ The mean-field term $\mathcal{S}_{\rm MF}=\hbar\beta\mathcal{A}\,\Omega_{\rm MF}$ is proportional to the the thermodynamic potential given in Eq. (S26) for the case of a symmetric mixture. Using the short-hand notation $(\mathbf{r},\tau)\equiv r$, $\int d\mathbf{r}\int_{0}^{\beta\hbar}d\tau\equiv\int dr$ and introducing the Nambu spinor
@@
$\mathcal{S}_{\rm B}=\frac{1}{2}\int dr\int dr^{\prime}\Psi^{\dagger}(r)\left[-\mathcal{G}^{-1}(r,r^{\prime})\right]\Psi(r^{\prime}).$ (S44)
In order to analyze the gaussian kernel, it is convenient to perform a Fourier transform and work in momentum space $(i\omega_{m},\mathbf{k})$, where $\omega_{m}=2\pi m/\beta$ ($m\in\mathbb{Z}$) is a bosonic Matsubara frequency. The results at zero temperature can then be obtained through the analytic continuation $\hbar\partial_{\tau}\rightarrow i\omega_{m}\rightarrow\omega+i0^{+}$. The (inverse) Green’s function of the system is given by
@@
where we have restricted to the symmetric mixtures of lower polaritons that we analyze in the main text. Notice that the poles of the Green’s function, given by the condition $\text{det}\left[\mathcal{G}^{-1}(\mathbf{k},\omega\rightarrow E_{\mathbf{k}})\right]=0$, coincide with the eigenvalues of the Bogoliubov matrix $\tilde{\mathcal{H}}(\mathbf{k})$ defined in Eq. (S23).
@@ V.2 Spatial density profile in the slab-geometry configuration
$\frac{\hbar^{2}}{8M_{*}}\frac{\left(n_{0}^{\prime}(x)\right)^{2}}{n_{0}(x)}+\Omega(n_{0}(x))=\text{const.}=0,$ (S64)
@@
    numerically solve the second order ODE (S63) using the initial conditions from i) in the interval $x_{*},+x_{*}]$, where $x_{*}>0$ is such that $n_{0}(-x_{*})=n_{0\rm D}$;